OpenStudy (anonymous):

For how many different integral values of b are both roots of x^2+bx-16=0? Please explain .

OpenStudy (anonymous):

discriminant =b^2-4ac=b^2-4*1*-16=b^2+64 it is always >=0 for any integral value of b roots are real for any real value.